# A Calculus of Inheritance

### v4

@@ Appendix A Expressive Asymmetry: Proofs
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the inheritance facilities of $\mathscr{L}_{1}$. ∎
Corollary A.4.
Inheritance-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
Appendix B Multi-Path Self-Reference in Other Systems
@@
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties in this semantics, so no route-priority disambiguation rule is introduced.

### v6

@@ Conclusion
We have $M\cong_{0}N$ but $M\not\cong_{1}N$, so ${\cong_{0}}\neq{(\cong_{1}|_{L_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), the lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$. ∎
Theorem C.8 (Expressive asymmetry).
In the language universe $\mathscr{U}$ (Definition C.3), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus: it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus can, but not conversely (Definition 3.17 of Felleisen (Felleisen, 1991)).
Proof.
We must show: (a) inheritance-calculus can macro-express every $\mathscr{U}$-construct that the lazy $\lambda$-calculus can macro-express, and (b) the converse fails.
Part (b): the lazy $\lambda$-calculus $\not\geq$ inheritance-calculus
The $\lambda$-calculus contains its own constructors ($\lambda$, application, variable) and can macro-express $\mathbf{let}$-binding (Lemma C.4), but cannot macro-express the inheritance constructors of $\mathscr{U}$ (Theorem C.7). Inheritance-calculus contains those constructors. Therefore the lazy $\lambda$-calculus is not at least as expressive as inheritance-calculus with respect to $\mathscr{U}$.
Part (a): inheritance-calculus $\geq$ the lazy $\lambda$-calculus
We show that inheritance-calculus can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express. The $\lambda$-calculus contains $\lambda$-abstraction, application, and variable; it can macro-express $\mathbf{let}$-binding (Lemma C.4). We must show that inheritance-calculus can macro-express all four of these constructs. The translation proceeds in two stages:
  1. (1)
    ANF conversion. The A-normal form transformation (Flanagan et al., 1993) replaces nested applications with $\mathbf{let}$-bindings; it is macro-expressible in the $\lambda$-calculus by Lemma C.5. In the reverse direction, $\mathbf{let}$-bindings are macro-expressible in the $\lambda$-calculus by Proposition 3.7 of Felleisen (Felleisen, 1991) (restated as Lemma C.4). Thus ANF conversion is a macro-expressible equivalence between the $\lambda$-calculus and its ANF fragment.
  2. (2)
    The translation $\mathcal{T}$. Each ANF $\lambda$-calculus construct maps to a fixed syntactic abstraction over inheritance-calculus (Theorem C.6), satisfying condition E4 of Felleisen (Felleisen, 1991, Definition 3.11).
By Lemma C.2, composing the two macro-expressible translations gives a macro-expressible translation from the $\lambda$-calculus (with $\mathbf{let}$) into inheritance-calculus. Since inheritance-calculus trivially contains its own constructors, it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express. ∎
Appendix D Multi-Path Self-Reference in Other Systems
@@
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Because attribute resolution aggregates results via set union (which is intrinsically idempotent), identical property definitions from the two routes merge without conflict. Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.